OpenStudy (anonymous):

HELP!!!!!! Find the solution to the system algebraically. Show all work for full credit. Leave your answers in fraction form. y=6x^2 7x+y=3

OpenStudy (phi):

the first equation says y is equal to 6x^2 use that definition of y in the 2nd equation: 7x + y = 3 replace y with 6x^2 7x + 6x^2 = 3 put into standard form: 6x^2 + 7x -3 =0 can you solve this quadratic ?

OpenStudy (anonymous):

I dont knoww how

OpenStudy (anonymous):

isn't there an equation ? to plug thisi n

OpenStudy (phi):

quadratic formula http://www.purplemath.com/modules/quadform.htm

OpenStudy (phi):

notice they want Leave your answers in fraction form

OpenStudy (anonymous):

how would your write that out